OpenStudy (anonymous):

solve x^4-15*x^2+56=0

OpenStudy (anonymous):

let y = x^2, then we have: \[y = x^2 \Rightarrow y^2 = x^4, x^4-15x^2+56 = 0 \iff y^2-15y+56 = 0\] We can factor that equation of y by thinking of two numbers that multiply to 56, but add to -15. Those numbers are -8 and -7: \[y^2-15y+56 \iff (y-8)(y-7) = 0\] Replacing y = x^2 gives: \[(x^2-8)(x^2-7) = 0\] So the four solutions are: \[x^2-7 = 0 \iff x = \pm \sqrt{7}, x^2-8 = 0 \iff x = \pm \sqrt{8}\]

OpenStudy (anonymous):

Y^2+15Y+56=0